Find the probability of drawing an ace and a jack when two cards are drawn (without replacement) from a standard deck of cards
Answer by longjonsilver(2297)

you don't say if either order is OK or an ace then a jack. I will do both:
1. ORDER matters:
P(ace THEN jack) = P(ace) + P(jack)

so, first, there are 4 aces and 52 cards ---> P(ace) = 4/52 = 1/13.
then there are 4 jacks and 51 cards --> P(jack) = 4/51

so, P(ace then jack) = 1/13 + 4/51 = 0.155 to 3dp

How about P(jack then ace)?..the same numbers apply, so P=0.155 too.

2. ORDER does not matter:
Now P(ace and a jack in any order) = P(ace then jack) or P(jack then ace), which is 2 times 0.155 --> 0.311 (use the full number in the calculation, not 0.155, as this is an approximate value).

jon.